Find the ph of a 0.3molar solution of h202 dissolved in 500ml of water
1
Expert's answer
2019-05-28T06:21:10-0400

Solution.

Hydrogen peroxide is a very weak acid, so we need the following formula to calculate pH.


"pH = \frac{pKa}{2} - \frac{lg(C)}{2}"

"C = \frac{n}{V}"

C = 0.6 M


"pH = \frac{11.58}{2} - \frac{lg(0.6)}{2}"

pH = 5.90

Answer:

pH = 5.90